# Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition

### v1

@@ Abstract
Vision-Language Models (VLMs) have demonstrated impressive capabilities in understanding visual content, but their reliability in safety-critical contexts remains under-explored. We introduce VERI (Visual Emergency Recognition Dataset), a carefully designed diagnostic benchmark of 200 images (100 contrastive pairs). Each emergency scene is matched with a visually similar but safe counterpart through multi-stage human verification and iterative refinement. Using a two-stage protocol—risk identification and emergency response—we evaluate 14 VLMs (2B-124B parameters) across medical emergencies, accidents, and natural disasters. Our analysis reveals a systematic overreaction problem: models excel at identifying real emergencies (70-100% success rate) but suffer from an alarming rate of false alarms, misidentifying 31-96% of safe situations as dangerous, with 10 scenarios failed by all models regardless of scale. This "better-safe-than-sorry" bias manifests primarily through contextual overinterpretation (88-93% of errors), challenging VLMs’ reliability for safety applications. These findings highlight persistent limitations that are not resolved by increasing model scale, motivating targeted approaches for improving contextual safety assessment in visually misleading scenarios.
[…]
tcb@breakable
Better Safe Than Sorry? Overreaction Problem of Vision Language Models in Visual Emergency Recognition
Dasol Choi1,2  Seunghyun Lee1  Youngsook Song3thanks: Corresponding Author Yonsei University1  MODULABS2  Lablup Inc.3 dasolchoi@yonsei.ac.kr  lutris@yonsei.ac.kr  yssong@lablup.com
@@ 1 Introduction
Vision-Language Models (VLMs) have made remarkable progress in understanding visual content, advancing from simple object recognition to sophisticated scene understanding and contextual reasoning (OpenAI, 2023; Google, 2023; Anthropic, 2024). These models now power a wide range of applications from content moderation to assistive technologies (Zhang et al., 2024). However, their reliability in safety-critical scenarios remains underexplored. This study investigates a critical question with real-world implications: To what extent can current VLMs distinguish between genuine emergencies and visually similar but safe scenarios?
False visual perceptions of emergencies can lead to costly resource mobilization. In October 2023, firefighters were dispatched to two separate incidents in New York and Seoul. In both cases, they discovered that high-definition fireplace videos playing on television screens had been mistaken for actual fires (Lloyd, 2023; Yoon, 2023). Such false alarms carry substantial costs. Each incident costs approximately \$1,000-2,500, with annual costs exceeding \$1 billion in the United States alone (Square, 2024). As VLMs power applications from smart home monitoring to CCTV surveillance, understanding their tendency to "overreact" becomes crucial (Yun and Lim, 2025).
Current research on safety-critical VLMs has primarily focused on specialized domains. These include medical diagnosis systems (Ben Abacha et al., 2019), autonomous driving perception (Caesar et al., 2020), and industrial robotics safety mechanisms (Dhillon et al., 2002). Our work addresses a distinct gap: everyday emergency recognition across domains. This task requires different contextual reasoning capabilities—distinguishing between visually similar scenarios based on subtle contextual cues rather than domain-specific features. Existing benchmarks typically evaluate models on isolated images, rarely using contrastive pairs that test the ability to distinguish between visually similar but semantically distinct situations.
To address these gaps, we introduce VERI (Visual Emergency Recognition Dataset) [^1] The VERI dataset is publicly available at
https://huggingface.co/datasets/Dasool/VERI-Emergency
, a diagnostic benchmark of 200 images comprising 100 contrastive pairs. Inspired by compact yet influential diagnostic benchmarks such as Winogender (720 examples) (Rudinger, 2018), WSC (60 examples) (Levesque et al., 2012), and PAIRS (200 images) (Fraser and Kiritchenko, 2024), VERI prioritizes diagnostic value over quantity. We employ controlled image pairs to evaluate VLMs’ ability to distinguish between emergency and non-emergency situations. Each pair underwent over 5 rounds of iterative refinement and verification to balance visual similarity with semantic distinction.
@@
  • Through evaluation of 14 VLMs (2B-124B parameters), we reveal a systematic "overreaction problem" persisting across architectures and scales, with 10 safe scenarios misclassified by all models.
@@
  • We show the overreaction problem persists despite model scaling, requiring targeted architectural improvements for contextual emergency assessment.
@@ 2 VERI: Visual Emergency Recognition Dataset
Effective emergency detection requires distinguishing between genuine threats and visually similar but safe situations. To evaluate this capability in Vision Language Models, we introduce VERI (Visual Emergency Recognition Dataset), a benchmark designed to challenge models’ contextual understanding in safety-critical scenarios.
@@ 2.1 Dataset Design and Taxonomy
The core design principle of VERI is the contrastive pair approach—each entry consists of two images with high visual similarity but fundamentally different semantic implications: one depicting a genuine emergency requiring intervention and the other showing a visually similar but safe scenario. As illustrated in Figure 2, these pairs include contrasts such as actual medical emergencies versus training simulations, genuine accidents versus staged scenarios, and real disasters versus controlled environments. This contrastive structure enables direct assessment of a model’s ability to distinguish between superficially similar scenes with different safety implications.
We organized VERI into three major categories of emergency situations (Accidents & Unsafe Behaviors, Personal Medical Emergencies, and Natural Disasters), as shown in Table 1. Each category presents unique visual challenges, from subtle physiological distress cues to complex environmental interpretation, as demonstrated in Figure 2.
@@ 2.2 Image Creation Process
Stage 1: Scenario Definition. For each category, we defined specific scenario pairs (e.g., "person requiring CPR" vs. "CPR training on mannequin"), identifying visual elements to preserve or alter for semantic distinction.
@@
Our process prioritized diagnostic value over quantity, focusing on expert-guided refinement to effectively challenge VLMs’ contextual reasoning.
@@ 2.3 Query Development
Q1: Risk Identification. The first query—"Does the situation in this image appear dangerous or safe?"—is a straightforward binary classification task designed to evaluate models’ ability to distinguish between dangerous and safe situations.
Q2: Emergency Response. For emergency images, we used the query—"Based on this situation, what emergency response should be taken?"—to assess models’ ability to recommend appropriate actions. We created a standardized response bank through a three-step process. First, we used image captions as input to Perplexity to generate initial responses based on authoritative sources. Next, three evaluators verified these responses against original sources. Finally, we compiled critical actions for each emergency type as evaluation criteria.
Statistic Count
Total image pairs 100
Total images 200
Accidents & Unsafe Behaviors pairs 35
Personal Medical Emergencies pairs 33
Natural Disasters pairs 32
Risk identification QA (Q1) 200
Emergency Response QA (Q2) 100
Detailed image captions 200
Table 2: VERI dataset statistics
@@ 2.4 Dataset Statistics
The final VERI dataset consists of 100 image pairs (200 total images) distributed across three categories: Accidents & Unsafe Behaviors (35 pairs), Personal Medical Emergencies (33 pairs), and Natural Disasters (32 pairs). For evaluation, we created 200 binary classification questions (Q1) covering all images and 100 open-ended response questions (Q2) for emergency images only. Each image is accompanied by a detailed caption describing the scene context, key elements, and situational details, providing additional textual information that can be used for multimodal training or analysis. Figure 2 illustrates representative examples from each category, demonstrating both the visual similarity within pairs and the semantic distinction between emergency and safe scenarios.
@@ 3.1 Models
We evaluated 14 open-source VLMs (2B-124B parameters) across different architectures: Qwen2.5-VL (Bai et al., 2025) (3B, 7B, 32B, 72B), transformer-based models optimized for visual recognition; LLaVA-Next (Liu et al., 2024) (7B, 13B), integrating CLIP encoders with large language models; InternVL3 (Zhu et al., 2025) (2B, 8B, 14B), using a "ViT-MLP-LLM" architecture; the Mistral family , including Mistral-Small (24B) (AI, 2025) and Pixtral variants (12B, 124B) (Agrawal et al., 2024); and other architectures represented by Idefics2 (8B) (Laurençon et al., 2024) and Phi-3.5-vision (4B) (Abdin et al., 2024). All models were evaluated using publicly available checkpoints without task-specific fine-tuning to assess zero-shot capabilities.
@@ 3.2 Evaluation Protocol
Q1 (Binary Classification). The model is asked to answer whether the situation in the image is dangerous or safe. It selects one of two choices (A. Dangerous / B. Safe) and provides a brief reasoning. We compute precision, recall, and F1 score based on human-annotated binary ground truth labels.
Q2 (Open-ended Response). For images classified as dangerous by the model and correctly aligned with the ground truth, we further ask what an appropriate emergency response should be. We evaluated these responses using GPT-4o as a judge, providing it with the image caption and our curated gold-standard answers as references. The judge was instructed to score each response on a scale from 0 to 1 based on its alignment with the reference materials and appropriateness for the emergency situation.
@@ 4.1 The Overreaction Phenomenon
To evaluate VLMs’ emergency recognition capabilities, we assessed their performance on both risk identification (Q1) and emergency response (Q2) tasks (Table 3). Our evaluation reveals a consistent pattern: models achieve high recall (0.70-1.00) in identifying dangerous situations, but precision is notably lower (0.51-0.72), indicating a systematic "better-safe-than-sorry" bias—an "overreaction problem." This pattern persists across model families and parameter scales (2B-124B). Model size does not consistently correlate with improved precision; InternVL3 (8B) achieves the highest precision (0.72), outperforming larger variants. False positive rates (safe images misclassified as dangerous) ranged from 31% to 96%—substantially higher than false negative rates (missed emergencies), which ranged from 0% to 30%. Figure 3(a) illustrates this bias through the precision-recall tradeoff, with models clustering in the high-recall but lower-precision region.
@@ 4.2 Emergency Response Evaluation
We found a disconnect between risk identification and response capabilities. InternVL3 (8B) achieved the highest precision in risk identification but only moderate response quality, while models with lower precision often produced better emergency responses. This suggests VLMs develop these capabilities through different mechanisms—procedural knowledge scaling more predictably than contextual reasoning. Even when models correctly identify emergencies, their ability to recommend appropriate interventions remains limited. Examples can be found in Appendix A.3.
@@ 4.3 Category-Specific Analysis
These findings suggest VLMs struggle most with scenarios where emergency status depends on fine-grained contextual details rather than obvious visual elements. For emergency response quality, Q1 showed a clearer category preference (ND > AB > PME with differences up to 0.07), while Q2 exhibited smaller average differences between categories (approximately 0.03). This suggests that once a model correctly identifies an emergency, its ability to generate appropriate responses is more consistent across different emergency types. For detailed analysis of model-specific performance across categories, see Appendix A.2.
@@ 4.4 Error Patterns and Analysis
Visual Misinterpretation: Models incorrectly perceive visual elements, failing to distinguish between safe and dangerous scenarios (e.g., confusing mannequins with real people, theatrical makeup with real injuries).
Contextual Overinterpretation: Models correctly identify visual elements but fail to properly interpret their safety implications within the broader context. This manifests as an exaggeration of potential risks by misweighing contextual factors. For example, models claim "the child’s shirt could lead to choking hazards" or "the car’s rearview mirror could blind the cyclist.
@@
These findings suggest that current VLMs can detect potentially hazardous elements, but lack the nuanced reasoning required to assess whether these elements pose actual dangers in specific contexts—a critical limitation for their reliability in safety applications. For detailed error pattern analysis across model sizes and categories, see Appendix C, and additional examples of Contextual Overinterpretation can be found in Appendix B.
@@ 4.5 Model Size and Category Effects
Our analysis reveals a non-linear relationship between model size and performance (Kaplan et al., 2020). The InternVL3 family exhibits a "Goldilocks effect," (Miller and Alexander, 2025) where the mid-sized 8B model excels in Accidents & Behaviors (0.83) and Natural Disasters (0.81), yet underperforms in Medical Emergencies (0.59) compared to both smaller 2B (0.75) and larger 14B (0.77) variants (see Table 4).
@@ 5.1 Balancing Safety and Accuracy
Our findings highlight a fundamental tension in visual emergency recognition: the trade-off between sensitivity to potential dangers and precision in distinguishing genuine emergencies from safe scenarios. While high sensitivity is desirable, the substantial false positive rates we observed (28-48%) could undermine practical utility in real-world applications. False alarms in emergency services lead to resource misallocation and "alarm fatigue," where users begin to distrust system warnings. Conversely, increasing precision must consider the severe consequences of missed emergencies. This optimization challenge varies across domains - medical triage might justifiably prioritize recall, while home monitoring systems require higher precision for user trust. The higher false positive rates in medical emergencies (38%) compared to natural disasters (27%) further suggest different domains require tailored approaches to balancing sensitivity and specificity.
@@ 5.2 Implications for Model Development
Our findings suggest several approaches for improving VLMs’ emergency recognition capabilities. Enhancing contextual reasoning is critical, as models struggle to weigh contextual factors when assessing dangers. The "Goldilocks effect"—where mid-sized models sometimes outperform larger ones—challenges the assumption that scaling parameters automatically improves safety performance. Category-specific optimization appears necessary; medical emergencies might benefit from specialized physiological feature detection, while disaster recognition could focus on environmental safety indicators. Contrastive learning with paired examples similar to our VERI dataset could help models better distinguish between visually similar but semantically distinct safety situations.

### v2

@@ Abstract
Vision-Language Models (VLMs) have shown capabilities in interpreting visual content, but their reliability in safety-critical everyday life scenarios remains insufficiently explored. We introduce VERI (Visual Emergency Recognition Dataset), a diagnostic benchmark comprising 200 images organized into 100 contrastive pairs. Each emergency scene is paired with a visually similar but safe counterpart through human verification and refinement. Using a two-stage evaluation protocol—risk identification and emergency response—we assess 14 VLMs (2B to 124B parameters) across medical emergencies, accidents, and natural disasters. Our analysis reveals an ”overreaction problem,” where models accurately identify genuine emergencies (70-100% success rate) but produce high false-positive rates, misclassifying 31-96% of safe situations as dangerous. Ten safe scenarios were universally misclassified by all models regardless of scale. This ”better-safe-than-sorry” bias primarily results from contextual overinterpretation (88-93% of errors), challenging VLM reliability in safety-critical applications. These findings highlight fundamental limitations in current VLM architectures, which persist despite increased model scale. Our results demonstrate an urgent need for strategies specifically improving contextual reasoning in ambiguous visual situations. The consistently low performance of the model indicates that these data serve effectively as a diagnostic data set.
@@ 1 Introduction
Vision-Language Models (VLMs) have made remarkable progress in understanding visual content, advancing from simple object recognition to sophisticated scene understanding and contextual reasoning [17, 10, 4]. These models now power a wide range of applications from content moderation to assistive technologies [26]. However, their reliability in safety-critical scenarios remains underexplored. This gap raises a critical question with real-world implications: To what extent can current VLMs distinguish between genuine emergencies and visually similar but safe scenarios? To investigate this question systematically, we utilized generated synthetic (fake) data, which allows us to create controlled comparisons between dangerous and benign situations while maintaining visual similarity.
False visual perceptions of emergencies can lead to costly resource mobilization. In October 2023, firefighters were dispatched to two separate incidents in New York and Seoul. In both cases, they discovered that high-definition fireplace videos playing on television screens had been mistaken for actual fires [15, 23]. Such false alarms carry substantial costs. Each incident costs approximately \$1,000-2,500, with annual costs exceeding \$1 billion in the United States alone [19]. As VLMs power applications from smart home monitoring to CCTV surveillance, understanding their tendency to ”overreact” becomes crucial [24]. As VLMs increasingly power applications from smart home monitoring to CCTV surveillance, understanding their tendency to ’overreact’ becomes crucial  [24]. Current research on safety-critical VLMs has primarily focused on specialized domains. These include medical diagnosis systems [6], autonomous driving perception [7], and industrial robotics safety mechanisms [8]. Our work addresses a distinct gap: everyday emergency recognition across domains. This task requires different contextual reasoning capabilities—distinguishing between visually similar scenarios based on subtle contextual cues rather than domain-specific features. Existing benchmarks typically evaluate models on isolated images, rarely using contrastive pairs that test the ability to distinguish between visually similar but semantically distinct situations.
To address these gaps, we introduce VERI (Visual Emergency Recognition Dataset), a diagnostic benchmark of 200 images comprising 100 contrastive pairs. Inspired by compact yet influential diagnostic benchmarks such as Winogender (720 examples) [18], WSC (60 examples) [13], and PAIRS (200 images) [9], VERI prioritizes diagnostic value over quantity. We employ controlled image pairs to evaluate VLMs’ ability to distinguish between emergency and non-emergency situations. Each pair underwent over 5 rounds of iterative refinement and verification to balance visual similarity with semantic distinction.
@@
  • Through evaluation of 14 VLMs (2B-124B parameters), we reveal a systematic ”overreaction problem” persisting across architectures and scales, with 10 safe scenarios misclassified by all models.
@@
  • We show that the overreaction problem persists despite model scaling, requiring targeted architectural improvements for contextual emergency assessment.
@@ 2.1 Dataset Design and Taxonomy
Effective emergency detection requires distinguishing between genuine threats and visually similar but safe situations. To address this challenge, the VERI dataset (200 images) has been specifically designed to study risks in everyday life, rather than focusing on specialized domains such as medical diagnosis or industrial safety. This focus on everyday risks is particularly relevant given the widespread adoption of autonomous driving and AI-driven robotics in our daily lives. By establishing this diagnostic benchmark, our study deliberately narrows its research scope to typical daily-life scenarios, where accurate threat detection is becoming increasingly critical. The core design principle of VERI is the contrastive pair approach—each entry consists of two images with high visual similarity but fundamentally different semantic implications: one depicting a genuine emergency requiring intervention and the other showing a visually similar but safe scenario. As illustrated in Figure 2, these pairs include contrasts such as actual medical emergencies versus training simulations, genuine accidents versus staged scenarios, and real disasters versus controlled environments. This contrastive structure enables direct assessment of a model’s ability to distinguish between superficially similar scenes with different safety implications. We organized VERI into three major categories of emergency situations (Accidents & Unsafe Behaviors, Personal Medical Emergencies, and Natural Disasters), as shown in Table 1. Each category presents unique visual challenges, from subtle physiological distress cues to complex environmental interpretation, as demonstrated in Figure 2.
@@ 2.2 Image Creation Process
Stage 1: Scenario Definition. For each category, we defined specific scenario pairs (e.g., ”person requiring CPR” vs. ”CPR training on mannequin”), identifying visual elements to preserve or alter for semantic distinction.
@@
Our process prioritized diagnostic value over quantity, focusing on expert-guided refinement to effectively challenge VLMs’ contextual reasoning. Because data for everyday emergency recognition did not previously exist, generating a large volume of realistic images by abstracting from existing raw data was impractical. Moreover, for diagnostic benchmarks with specific evaluation goals, the qualitative value and the control of variables are significantly more important than sheer quantity. This principle aligns with other influential diagnostic benchmarks in the literature. For instance, the PAIRS benchmark comprises 200 images, and the study ’Judging LLM-as-a-Judge with MT-Bench and Chatbot Arena’ (Zheng et al., 2023) evaluated models using only 80 carefully constructed questions across eight fields.
@@ 2.3 Query Development
Q1: Risk Identification. The first query—”Does the situation in this image appear dangerous or safe?”—is a straightforward binary classification task designed to evaluate models’ ability to distinguish between dangerous and safe situations.
Q2: Emergency Response. For emergency images, we used the query—”Based on this situation, what emergency response should be taken?”—to assess models’ ability to recommend appropriate actions. We created a standardized response bank through a three-step process. First, we used image captions as input to Perplexity to generate initial responses based on authoritative sources. Next, three evaluators verified these responses against original sources. Finally, we compiled critical actions for each emergency type as evaluation criteria.
@@ 2.4 Dataset Statistics
The final VERI dataset consists of 100 image pairs (200 total images) distributed across three categories: Accidents & Unsafe Behaviors (35 pairs), Personal Medical Emergencies (33 pairs), and Natural Disasters (32 pairs), as summarized in Table 3. For evaluation, we created 200 binary classification questions (Q1) covering all images and 100 open-ended response questions (Q2) for emergency images only. Each image is accompanied by a detailed caption describing the scene context, key elements, and situational details, providing additional textual information that can be used for multimodal training or analysis. Figure 2 illustrates representative examples from each category, demonstrating both the visual similarity within pairs and the semantic distinction between emergency and safe scenarios.
Statistic Count
Total image pairs 100
Total images 200
Accidents & Unsafe Behaviors pairs 35
Personal Medical Emergencies pairs 33
Natural Disasters pairs 32
Risk identification QA (Q1) 200
Emergency Response QA (Q2) 100
Detailed image captions 200
Table 3: VERI dataset statistics
@@ 3.1 Models
We evaluated 14 open-source VLMs (2B-124B parameters) across different architectures: Qwen2.5-VL [5] (3B, 7B, 32B, 72B), transformer-based models optimized for visual recognition; LLaVA-Next [14] (7B, 13B), integrating CLIP encoders with large language models; InternVL3 [27] (2B, 8B, 14B), using a ”ViT-MLP-LLM” architecture; the Mistral family , including Mistral-Small (24B) [3] and Pixtral variants (12B, 124B) [2]; and other architectures represented by Idefics2 (8B) [12] and Phi-3.5-vision (4B) [1]. All models were evaluated using publicly available checkpoints without task-specific fine-tuning to assess zero-shot capabilities.
@@ 3.2 Evaluation Protocol
Q1 (Binary Classification).
The model is asked whether the situation in the image is dangerous or safe. It selects one of two choices (A. Dangerous / B. Safe) and provides a brief reasoning. We compute precision, recall, and F1 score based on human-annotated binary ground truth labels.
Q2 (Open-ended Response).
For images classified as dangerous by the model and correctly aligned with the ground truth, we further ask what an appropriate emergency response should be. We evaluated these responses using GPT-4o as a judge, providing it with the image caption and our curated gold-standard answers as references. The judge was instructed to score each response on a scale from 0 to 1 based on its alignment with the reference materials and appropriateness for the emergency situation.
@@ 4.1 The Overreaction Phenomenon
To evaluate VLMs’ emergency recognition capabilities, we assessed their performance on both risk identification (Q1) and emergency response (Q2) tasks (Table 2). Our evaluation reveals a consistent pattern: models achieve high recall (0.70-1.00) in identifying dangerous situations, but precision is notably lower (0.51-0.72), indicating a systematic ”better-safe-than-sorry” bias—an ”overreaction problem.” This pattern persists across model families and parameter scales (2B-124B). Model size does not consistently correlate with improved precision; InternVL3 (8B) achieves the highest precision (0.72), outperforming larger variants. False positive rates (safe images misclassified as dangerous) ranged from 31% to 96%—substantially higher than false negative rates (missed emergencies), which ranged from 0% to 30%. Figure 3(a) illustrates this bias through the precision-recall tradeoff, with models clustering in the high-recall but lower-precision region.
@@ 4.2 Emergency Response Evaluation
We found a disconnect between risk identification and response capabilities. InternVL3 (8B) achieved the highest precision in risk identification but only moderate response quality, while models with lower precision often produced better emergency responses. This suggests VLMs develop these capabilities through different mechanisms—procedural knowledge scaling more predictably than contextual reasoning. Even when models correctly identify emergencies, their ability to recommend appropriate interventions remains limited. Examples can be found in the supplementary material.
@@ 4.3 Category-Specific Analysis
These findings suggest VLMs struggle most with scenarios where emergency status depends on fine-grained contextual details rather than obvious visual elements. For emergency response quality, Q1 showed a clearer category preference (ND $>$ AB $>$ PME with differences up to 0.07), while Q2 exhibited smaller average differences between categories (approximately 0.03). This suggests that once a model correctly identifies an emergency, its ability to generate appropriate responses is more consistent across different emergency types. For detailed model-specific performance breakdowns across emergency categories, see the supplementary material.
@@ 4.4 Error Patterns and Analysis
Visual Misinterpretation
Models incorrectly perceive visual elements, failing to distinguish between safe and dangerous scenarios (e.g., confusing mannequins with real people, theatrical makeup with real injuries).
Contextual Overinterpretation
Models correctly identify visual elements but fail to properly interpret their safety implications within the broader context. This manifests as an exaggeration of potential risks by misweighing contextual factors. For example, models claim “the child’s shirt could lead to choking hazards” or ”the car’s rearview mirror could blind the cyclist.
@@
These findings suggest that current VLMs can detect potentially hazardous elements, but lack the nuanced reasoning required to assess whether these elements pose actual dangers in specific contexts—a critical limitation for their reliability in safety applications. Detailed error pattern analysis across model sizes and categories, along with additional examples of Contextual Overinterpretation, are provided in the supplementary material.
@@ 4.5 Model Size and Category Effects
Our analysis reveals a non-linear relationship between model size and performance [11]. The InternVL3 family exhibits a “Goldilocks effect,” [16] where the mid-sized 8B model excels in Accidents & Behaviors (0.83) and Natural Disasters (0.81), yet underperforms in Medical Emergencies (0.59) compared to both smaller 2B (0.75) and larger 14B (0.77) variants (see Table 4).
@@ 5.1 Balancing Safety and Accuracy
Our findings highlight a fundamental tension in visual emergency recognition: the trade-off between sensitivity to potential dangers and precision in distinguishing genuine emergencies from safe scenarios. While high sensitivity is desirable, the substantial false positive rates we observed (28-48%) could undermine practical utility in real-world applications. False alarms in emergency services lead to resource misallocation and ”alarm fatigue,” where users begin to distrust system warnings. Conversely, increasing precision must consider the severe consequences of missed emergencies. This optimization challenge varies across domains - medical triage might justifiably prioritize recall, while home monitoring systems require higher precision for user trust. The higher false positive rates in medical emergencies (38%) compared to natural disasters (27%) further suggest that different domains require customized approaches to balance sensitivity and specificity.
@@ 5.2 Implications for Model Development
Our findings highlight potential directions for improving VLMs’ emergency recognition capabilities. Enhancing contextual reasoning remains crucial, as models currently struggle to appropriately interpret contextual nuances. We observed that mid-sized models can sometimes achieve performance comparable to or even slightly better than larger models, indicating that simply increasing model size may not always yield significant improvements. However, given our relatively small sample size and modest performance differences observed between model families, further investigation with larger and more diverse datasets is needed. Additionally, category-specific optimizations and contrastive learning approaches could enhance models’ differentiation abilities.
5.3 Limitations
Our binary classification methodology inherently limits the exploration of nuanced severity gradations within emergency scenarios, which are crucial for realistic emergency assessments. Moreover, although synthetic images allow precise control over visual elements essential for diagnostic purposes, they may include subtle artifacts not representative of the complexities present in real-world photographs, potentially affecting generalization.
Our evaluation protocol simplifies the inherently complex nature of real-world emergency decision-making, and judgments provided by GPT-4o, despite their consistency, may not always fully align with expert evaluations. The VERI dataset, comprising 200 images structured into 100 contrastive pairs, intentionally emphasizes diagnostic depth over breadth. Consequently, this focused design inevitably poses limitations regarding broader generalizability, particularly across diverse cultural and geographical contexts.
Controlled laboratory evaluations using curated image pairs may not entirely predict model performance under realistic deployment conditions, characterized by imperfect inputs and variability. However, despite these acknowledged limitations, the consistently low performance scores observed in the evaluated models underscore the diagnostic value of this dataset. Future research should seek to mitigate these limitations through larger-scale datasets, more detailed classification schemas, and validations conducted under less controlled, more realistic scenarios.